OpenStudy (anonymous):

prove the following identity: (sin^3 y -1)/(sin y -1) = sin^2 y + sin y + 1

OpenStudy (anonymous):

this has nothing whatsoever to do with trig. use \[\frac{a^3-1}{a-1}=\frac{(a-1)(a^2+a+1}{a-1}=a^2+a+1\] replace \[a\] by \[\sin(y)\] so get your answer

myininaya (myininaya):

ha you forgot to put a closing parenthesis

OpenStudy (anonymous):

grrr